If the perimeter of a rectangle is 24 ft and the width is 7 ft, find its length. [Hint: The perimeter of a rectangle is given by P = 2L + 2W.]
Answer by checkley75(3666)

2*7+2x=24
14+2x=24
2x=24-14
2x=10
x=10/2
x=5 ft. answer for the length.
proof
2*7+2*5=24
14=10=24
24=24